Hi I am a parent trying to help my son in Algebra homework...
We understand some of it but not from the second line down. Your help is really needed..
7(6-y)=-3(y-2)
42-7y = -3y-6 {what number do we add or subtract to each side & why}
42-7y+3y = -3y-6+3y
42+y = -6
42+y-42 = -6-42
y=48
Is THAT CORRECT?

Answer by Here2Help(19)

Here you go:
7(6-y)=-3(y-2)
42-7y=-3y+6 (You should add 7 to both sides so you can have a positive #)
42+7y-7y=+7y-3y+6
42=4y+6 (+7y-7y=0 and cancel each other out)
42-6=4y+6-6 (+6-6=0 and cancel each other out)
36=4y
36/4=4y/4
9=y